# DemoRank: Selecting Effective Demonstrations for Large Language Models in Ranking Task

### v1

@@ 4.2 Demonstration Retriever DRetriever
In this part, we train DRetriever to retrieve potentially useful demonstrations for subsequent demonstration reranking. We apply an LLM to score a set of demonstration candidates to obtain supervised signals and use them to train the retriever through a multi-task learning strategy.
Scoring with LLM
@@
For a training input $I=(q,p)$ which contains a query-passage pair, we select a set of demonstrations from demonstration pool $\mathcal{P}$ as training candidates. Following previous studies Wang et al. (2023), we employ the BM25 algorithm to retrieve top-$b$ demonstrations. Due to the complex nature of passage ranking, the utility of a demonstration is not directly related to its similarity to the input Drozdov et al. (2023). To include more potential useful demonstrations for training, we also randomly sample another $B$ demonstrations from $\mathcal{P}$. The total number of training candidates is annotated as $N$ ($N=2*b$).
@@
$f(z_{i},I)=\frac{\text{Pr}(y|T,z_{i},I)}{\sum_{y^{\prime}\in Y}\text{Pr}(y^{\prime}|T,z_{i},I)},$ (3)
where $y$ is the relevance label for the query-passage pair in $I$, $Y=\{``\text{Yes}",``\text{No}"\}$ is the label space and $T$ is the task description. In this paper, the LLM scorer uses the same model as the LLM passage ranker. Nevertheless, we also explored the transferability of LLM scorer on different LLM passage rankers (see Appendix C).
Training
@@
$S(I,z_{i})=E_{I}(I)^{\top}E_{z}(z_{i}),$ (4)
where the two encoders $E_{I}$ and $E_{z}$ share parameters and encode with average pooling.
@@
$L_{\text{c}}=-\log\frac{e^{S(I,z^{+})}}{\sum_{z^{\prime}\in Z}e^{S(I,z^{\prime})}},$ (5)
where $Z=\{z^{+},z_{1}^{-},\dots,z_{N-1}^{-}\}$. Here we choose not to use in-batch negatives. The reasons are discussed in appendix D.
To make use of the fine-grained supervision of LLM’s feedback, we also consider a ranking loss RankNet Burges et al. (2005) to inject the ranking signal of candidates into training:
$L_{\text{r}}=\sum_{i,j}^{|Z|}\mathbb{1}_{r_{i}<r_{j}}*\log(1+e^{S(I,z_{j})-S(I,z_{i})}),$ (6)
@@
The final loss function $L$ is defined as the weighted sum of $L_{c}$ and $L_{r}$:
$L=\lambda L_{c}+L_{r},$ (7)
@@
4.3 Demonstration Reranker DReranker
Previous studies Wang et al. (2023); Rubin et al. (2022); Li et al. (2023) mainly use the top-$k$ retrieved demonstrations for ICL which ignores the demonstration dependencies and could be sub-optimal for ranking tasks. To mitigate this issue, we formulate the selection of the optimal $k$-shot permutation from retrieved demonstrations into a demonstration reranking problem and construct a novel dependency-aware training samples in an efficient way for the reranker’s training.
@@ 4.3 Demonstration Reranker DReranker
0:  Dependency-aware training samples $Y$.
1:  $Y\leftarrow\{\}$, selected demonstrations $S\leftarrow[]$.
2:  Retrieve top-$M$ demonstrations $Z^{\text{r}}$
3:  for $y=K$ to $1$ do
4:     // $y$ is the current ranking label.
5:     $z^{*}=\operatorname{arg}\max_{z_{j}\in Z^{\text{r}}\backslash S}f(\left[S,z_{j}\right],I)$, using Equation (3)
6:     $S\leftarrow\left[S,z^{*}\right]$, $Y\leftarrow Y\cup\{(z^{*},y)\}$
7:  end for
8:  for $z_{j}$ in $Z^{\text{r}}\backslash S$ do
9:     $Y\leftarrow Y\cup\{(z_{j},0)\}$
10:  end for
11:  return  $Y$
Constructing Dependency-aware Training Samples.
To align with the aim of our DReranker, we propose constructing a dependency-aware training samples for training. Specifically, given a training input $I$, we use our trained DRetriever to retrieve top-$M$ demonstrations $Z^{\text{r}}$ from the demonstration pool. Then, we iteratively select demonstrations from $Z^{\text{r}}$ and annotate each of them with a ranking label, as Figure 2 shows. In each iteration, we select, from the unselected demonstrations in $Z^{\text{r}}$, the one that maximizes the LLM’s feedback when concatenated with already selected ones. Once a demonstration is selected, we append it to the training samples. This process considers previous demonstration sequence when selecting the current demonstration and approximates the optimal $k$-shot demonstration permutation incrementally, which is time-efficient and aligns with the few-shot setting. Note that as the number of iterations increases, the computational cost of LLM inference also increases. Due to limited computational resources, we set a maximum iteration number $K$. After the $K$-th iteration is completed, we annotate a ranking label from $K$ to 1 to each demonstration in the training sample according to their selection order and annotate 0 to the unselected demonstrations in $Z^{\text{r}}$. Algorithm 1 shows this procedure.
Training
After constructing the dependency-aware training sample, we obtain a ranking label for each demonstration candidate in $Z^{\text{r}}$. We employ a cross-encoder model to train our DReranker. The model takes as input the concatenation of training input $I$ and one candidate $z_{i}$ with a “[SEP]” token and outputs a prediction score $s_{i}$ using the representation of “[CLS]” token. Then we apply the RankNet loss function to optimize the reranker model, similar to Equation (6):
$L_{\text{r}}=\sum_{i,j}^{|Z^{\text{r}}|}\mathbb{1}_{y_{i}>y_{j}}*\log(1+e^{s_{j}-s_{i}}),$ (8)
where $y_{i}$ represents the ranking label of $z_{i}$. Note that our DReranker only receives an input and a single demonstration, without including dependent demonstrations, which may not fully capture the dependency-aware ranking labels. Nonetheless, this design saves inference time, making our DReranker more efficient. We plan to explore architectures that can model multiple dependent demonstrations efficiently in the future.
4.4 Inference
@@
During inference, we first encode the entire demonstration pool $\mathcal{P}$ using our trained DRetriever and build the index. Then, given a test input $I^{\text{test}}=(q^{\text{test}},p^{\text{test}}_{i})$, we retrieve top-$M$ demonstrations using DRetriever and rerank them using our trained DReranker. Finally, we choose top-$k$ reranked demonstrations as the in-context demonstrations and concatenate them with the test input to calculate the relevance score. We perform this process for all retrieved passages of $q^{\text{test}}$ and rank these passages based on their relevance scores.
5 Experiments
5.1 Setting
Datasets
@@ 5.1 Setting
In our experiments, we train and evaluate our DemoRank on diverse ranking datasets, including HotpotQA Yang et al. (2018), NQ Kwiatkowski et al. (2019), FEVER Thorne et al. (2018) and MS MARCO Nguyen et al. (2016). We use their training set to train our models respectively and evaluate the models on the corresponding test set (for MS MARCO, the evaluation is conducted on its development set as well as two in-domain datasets, TREC DL19 Craswell et al. (2020b) and TREC DL20 Craswell et al. (2020a)).
Implementation Details
We use FLAN-T5-XL Chung et al. (2022) as the frozen LLM for demonstration scoring and passage ranking unless otherwise specified. During the training stage, the number of demonstration candidates for retriever and reranker ($N$ and $M$ respectively) are both set as 50. And the maximum iteration number $K$ in Section 4.3 is set as 4. During training, we apply e5-base-v2 Wang et al. (2022) and DeBERTa-v3-base He et al. (2023) to initialize our demonstration retriever and reranker respectively. Following previous studies Sun et al. (2023); Zhuang et al. (2023a), we use the top-100 passages retrieved by BM25 as the passages to rerank. Due to the limited space, more implementation details on model training and inference are listed in Appendix A.
Baselines
$\bullet$ DBS Drozdov et al. (2023): DBS is a rule-based selection approach based on query generation in passage ranking. It selects the demonstrations which are the most difficult for the LLM to predict. In this paper, we implemented the algorithm based on the relevance generation approach. We define a score for each demonstration as the probability of the LLM generating the corresponding relevance label given a query and passage. The demonstrations with the lowest scores are applied.
$\bullet$ K-means: K-means is another static demonstration selection approach. This method clusters all the demonstrations in the pool into $k$ clusters and then selects $k$ demonstrations closest to each cluster center for ICL. We use the E5 Wang et al. (2022) model to obtain the demonstration embeddings for clustering.
$\bullet$ BM25 Robertson and Zaragoza (2009): BM25 is a widely-used sparse retriever. We apply BM25 to retrieve demonstrations that are most similar to the test query.
$\bullet$ SBERT Reimers and Gurevych (2019): We use Sentence-BERT as the off-the-shelf demonstration retriever following Rubin et al. (2022) [^1] The checkpoint is from https://huggingface.co/sentence-transformers/paraphrase-mpnet-base-v2.. We use SBERT to encode all the demonstrations in the pool and retrieve the most similar demonstrations.
$\bullet$ E5 Wang et al. (2022): E5 is another off-the-shelf dense retriever. Following Wang et al. (2023), we use the same retrieval method as SBERT based on e5-base-v2 checkpoint [^2] https://huggingface.co/intfloat/e5-base-v2.
Method HotpotQA NQ FEVER DL19 DL20 MS MARCO Avg
@@
1-shot Random 59.71 48.69 38.41 66.76 65.35 33.53 52.08
K-means 59.62 48.68 37.96 66.45 65.30 33.59 51.93
DBS 60.34 49.05 38.96 66.83 65.79 33.54 52.42
BM25 61.46 49.53 40.43 65.08 65.86 33.73 52.68
SBERT 58.41 49.49 36.25 66.63 64.18 33.98 51.49
E5 61.70 49.49 39.96 66.48 65.20 33.79 52.77
DemoRank 65.64 52.11 44.16 68.64 67.38 35.03 55.49
3-shot Random 59.42 48.61 38.61 66.57 64.84 33.70 51.96
@@
DemoRank 66.39 52.52 46.90 68.28 67.66 35.12 56.15
Table 1: Main results (NDCG@10) on different datasets. The best results are marked in bold and the column Avg represents the average performance of all datasets. The Initial Order represents the order of the top-100 passages retrieved by BM25.
5.2 Main Results
We compare DemoRank with baselines in 1-shot and 3-shot ICL respectively. Note that although DemoRank mainly focuses on few-shot scenarios, it can also work in 1-shot ICL, so we provide the performance of the 1-shot ICL as a reference. Table 1 shows the main results of our experiments. From the results, we draw the following observations: (1) Our framework DemoRank outperforms all the baselines significantly across all datasets. For example, in 3-shot ICL, DemoRank outperforms the second-best model E5 on HotpotQA by 3 points, and the second-best model BM25 on FEVER by about 7 points. It shows the DemoRank’s powerful ability to select demonstrations. (2) When expanding from 1-shot to 3-shot, DemoRank shows greater improvement on Avg metric compared to other baselines, indicating that our DemoRank can better enhance the ICL performance in few-shot scenario. (3) The similarity-based demonstration selection baselines (e.g., E5) outperform Random, K-means, and DBS baselines, but still lags far behind DemoRank, which proves the effectiveness of task-specific finetuning based on LLM’s feedback.
Method NQ DL19 FEVER Avg

### v2

@@ 4.2. Demonstration Retriever DRetriever
In this part, we train DRetriever to retrieve potentially useful demonstrations for subsequent dependency-aware reranking. We apply an LLM to score a set of demonstration candidates to obtain supervised signals and use them to train the retriever through a multi-task learning strategy.
4.2.1. Scoring with LLM
@@
For a training input $I=(q,p)$ which contains a query-passage pair, we select a set of demonstrations from demonstration pool $\mathcal{P}$ as training candidates. Following previous studies (Wang et al., 2023), we employ the BM25 algorithm to retrieve top-$b$ demonstrations. Due to the complex nature of passage ranking, the utility of a demonstration is not directly related to its similarity to the input (Drozdov et al., 2023). To include more potential useful demonstrations for training, we also randomly sample another $b$ demonstrations from $\mathcal{P}$. The total number of training candidates is annotated as $N$ ($N=2*b$).
@@
(3) $f(z_{i},I)=\frac{\text{Pr}(y|T,z_{i},I)}{\sum_{y^{\prime}\in Y}\text{Pr}(y^{\prime}|T,z_{i},I)},$
where $y$ is the relevance label for the query-passage pair in $I$, $Y=\{``\text{Yes}",``\text{No}"\}$ is the label space and $T$ is the task description. In this paper, the LLM scorer uses the same model as the LLM passage ranker. Nevertheless, we also explored the transferability of LLM scorer on different LLM passage rankers in Section 5.3.5).
4.2.2. Training
@@
(4) $S(I,z_{i})=E_{I}(I)^{\top}E_{z}(z_{i}),$
where the two encoders $E_{I}$ and $E_{z}$ share same parameters and encode with average pooling.
@@
(5) $L_{\text{c}}=-\log\frac{e^{S(I,z^{+})}}{\sum_{z^{\prime}\in Z}e^{S(I,z^{\prime})}},$
where $Z=\{z^{+},z_{1}^{-},\dots,z_{N-1}^{-}\}$. Here we choose not to use in-batch negatives. The reasons are discussed in Section 5.3.1.
To make use of the fine-grained supervision of LLM’s feedback, we also consider a ranking loss RankNet (Burges et al., 2005) to inject the ranking signal of candidates into training:
(6) $L_{\text{r}}=\sum_{i,j}^{|Z|}\mathds{1}_{r_{i}<r_{j}}\log(1+e^{S(I,z_{j})-S(I,z_{i})}),$
@@
The final loss function $L$ is defined as the weighted sum of contrastive loss $L_{c}$ and ranking loss $L_{r}$:
(7) $L=\lambda L_{c}+L_{r},$
@@
4.3. Demonstration Reranker DReranker
Previous studies (Wang et al., 2023; Rubin et al., 2022; Li et al., 2023) mainly use the top-$k$ retrieved demonstrations for ICL which ignores the demonstration dependencies and could be sub-optimal for ranking tasks. In this part, we introduce a novel demonstration reranker DReranker to rerank the retrieved demonstrations to make the top-$k$ more suitable for $k$-shot ICL. As mentioned in Section 1, generating training data for the DReranker presents significant challenges: (1) LLM’s feedback on individual demonstrations used for retriever training fails to capture the demonstration dependencies; (2) constructing the ground truth ranking for $k$-shot demonstration requires using LLM scorer to evaluate all possible permutations, which has high time complexity and impossible. To overcome these challenges, we design a time-efficient approach to iteratively approximate the optimal demonstration sequence and generate a series of scored demonstration lists at each iteration, which will be used as the training samples for DReranker. In this process, the LLM scorer evaluates multiple demonstrations at a time, thus taking the demonstration dependencies into account. Finally, we propose a list-pairwise training approach which compares a pair of demonstration lists (with only the last demonstration being different). This approach teaches the DReranker how to select the next demonstration given the previously selected ones.
@@ 4.3. Demonstration Reranker DReranker
0:  Dependency-aware training samples $O$.
1:  Retrieve top-$M$ demonstrations $Z^{\text{r}}$
2:  $O\leftarrow\{\}$, selected demonstrations $Z^{s}\leftarrow[]$, unselected demonstrations $Z^{u}\leftarrow Z^{\text{r}}$
3:  for $i=1$ to $K$ do
4:     Score and rank each $z_{j}\in Z^{u}$ using $f(Z^{s}\oplus z_{j},I)$ based on Equation (3)
5:     Append $(I,Z^{s},Z^{u})$ and the ranking of scored $Z^{u}$ to $O$
6:     Select a demonstration $z^{*}\in Z^{u}$ based on Equation (8)
7:     Append $z^{*}$ to $Z^{s}$, remove $z^{*}$ from $Z^{u}$
8:  end for
9:  return  $O$
4.3.1. Constructing Dependency-aware Training Samples.
As we mentioned above, our Dreranker acts as a sequential demonstration selector. To generate its training data, we propose to approximate the optimal demonstration sequence iteratively and construct a kind of dependency-aware training samples. Specifically, given a training input $I$, we use our trained DRetriever to retrieve top-$M$ demonstrations $Z^{\text{r}}$ from the demonstration pool. For the first iteration, we use the LLM scorer to score and rank these demonstrations based on Equation (3). Then, we design a sampling strategy to select a demonstration from the unselected demonstrations $Z^{\text{u}}$ ($Z^{\text{u}}$=$Z^{\text{r}}$ for the first iteration) based on its rank:
(8) $p(z_{i})=\frac{g(z_{i})}{\sum_{z_{j}\in Z^{\text{u}}}g(z_{j})},\quad g(z_{i})=\text{exp}(-r_{i}),$
where $r_{i}$ is the rank of demonstration $z_{i}$. This strategy ensures that higher-scoring demonstrations are more likely to be selected for constructing the optimal demonstration sequence. Meanwhile, it retains the possibility of selecting lower-scoring demonstrations, which helps DReranker select the next demonstration when the previous sequence is not good enough, thus improving its ability to handle previous sequences of different quality. For the $k$-th iteration, we use the LLM scorer to score each unselected demonstration by concatenating it with previously selected demonstration sequence and obtain a ranked list. Then we select a demonstration based on the same sampling strategy. Meanwhile, we obtain a ranking of $k$-shot demonstration list (e.g., 3-shot ranking $([z_{3},z_{1},z_{2}]>[z_{3},z_{1},z_{4}]>\dots)$ in Figure 2), which acts as a dependency-aware training sample and will be used by our list-pairwise training approach in the next section. In this way, the dependency between each demonstration and previous demonstrations is considered. Note that as the number of iterations increases, the computational cost of LLM inference also increases. Due to limited computational resources, we set a maximum iteration number $K$. Algorithm 1 shows the whole constructing process.
4.3.2. List-pairwise Training
Learning to rank approaches are usually divided into three categories: pointwise, pairwise and listwise (Liu et al., 2009). Demonstration reranking is naturally a listwise problem because the score of a demonstration depends on previous demonstrations. Take the 3-shot demonstration in Figure 2 as an example, given that $z_{3}$ and $z_{1}$ have been selected, $z_{2}$ have a higher LLM score than $z_{4}$. In this part, we propose a list-pairwise training method to optimize our DReranker. We call it list-pairwise because the loss is calculated by comparing a pair of demonstration lists ($l_{1},l_{2}$) where $l_{1}$ and $l_{2}$ differ only in the last demonstration. With the dependency-aware training samples from 1-shot to $K$-shot already constructed, we generate all pairs of demonstration lists for each shot (for 1-shot, the list length is 1). Then, we calculate our list-pairwise loss similarly to the pairwise RankNet loss:
(9) $L_{\text{list-pairwise}}=\sum_{k=1}^{K}\sum_{i,j}^{|L_{k}|}\mathds{1}_{r_{i}<r_{j}}\log\left(1+e^{\text{score}(I,l_{j})-\text{score}(I,l_{i})}\right),$
where $L_{k}$ is the ranked list of $k$-shot demonstration list, $r_{i}$ is the rank of $i$-th demonstration list $l_{i}$ in $L_{k}$. Our DReranker is based on a cross-encoder model, which calculates $\text{score}(I,l_{i})$ by taking the concatenation of the training input $I$ and demonstration list $l_{i}$ as input, and output a score using the representation of “[CLS]” token. Observing demonstration lists of different shots helps the DReranker learn to select the optimal demonstration under a given previous demonstration sequence.
4.4. Inference
@@
During inference, we first encode the entire demonstration pool $\mathcal{P}$ using our trained DRetriever and build the index. Then, given a test input $I^{\text{test}}=(q^{\text{test}},p^{\text{test}}_{i})$, we first retrieve top-$D$ demonstrations using DRetriever. Then, we sequentially and greedily select the highest-scoring demonstration and append it to the ranking list. Specifically, the first demonstration is selected only conditioned on the test input $I^{\text{test}}$. Once the top-$i$ demonstrations have been selected, we choose the demonstration from the unselected ones that, when concatenated with the selected top-$i$ demonstration and $I^{\text{test}}$, results in the highest LLM score, and append it to the ranking list. This process is repeated until the ranking list contains $K$ demonstrations. Finally, we choose these $k$ demonstrations as the in-context demonstrations and concatenate them with the test input to calculate the relevance score. We perform this process for all retrieved passages of $q^{\text{test}}$ and rank these passages based on their relevance scores.
5. Experiments
5.1. Setting
5.1.1. Datasets
@@
In our experiments, we train and evaluate our DemoRank on diverse ranking datasets, including HotpotQA (Yang et al., 2018), NQ (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018) and MS MARCO (Nguyen et al., 2016). We use their training set to train our models respectively and evaluate the models on the corresponding test set (for MS MARCO, the evaluation is conducted on its development set as well as two in-domain datasets, TREC DL19 (Craswell et al., 2020b) and TREC DL20 (Craswell et al., 2020a)).
5.1.2. Implementation Details
We use FLAN-T5-XL (Chung et al., 2022) as the LLM scorer and passage ranker unless otherwise specified. For the construction of training input, we pair each query in the training set with one relevant passage and one irrelevant passage respectively, thus generating two training inputs. Next, we will introduce the training details of our models. As for DRetriever, we set the number of demonstration candidates $N$ as 50 and hyper-parameter $\lambda$ for multi-task learning as 0.2. Following previous study (Wang et al., 2023), we use e5-base-v2 (Wang et al., 2022) to initialize DRetriever. As for DReranker, the number $M$ of retrieved demonstrations for scoring is 50 and the maximum iteration number $K$ in Section 4.3 is set as 3. We apply DeBERTa-v3-base (He et al., 2023) for model initialization. Both the DRetriever and DReranker are trained for 2 epochs. During inference, the number of retrieved demonstration $D$ for reranking is set as 30.
Table 1. Main results (NDCG@10) on different datasets. The best results are marked in bold and the column Average represents the average performance of all datasets. $\dagger$” indicates the model outperforms the best baseline significantly with paired t-test at $p$-value $<$ 0.05 level.
Method HotpotQA NQ FEVER DL19 DL20 MS MARCO Average
@@ 5.1.2. Implementation Details
Random 59.42 48.61 38.61 66.57 64.84 33.70 51.96
@@
DemoRank 67.03 52.31 48.43 67.76 66.76 35.31 56.27
5.1.3. Baselines
$\bullet$ Initial Order: Following previous studies (Sun et al., 2023; Zhuang et al., 2023b), we use the top-100 passages retrieved by BM25 as the initial passages order. All the subsequent baselines will perform reranking on these passages.
$\bullet$ 0-shot: The passage ranking approach based on relevance generation without any demonstration.
$\bullet$ DBS (Drozdov et al., 2023): DBS is a rule-based selection approach based on query generation in passage ranking. It selects the demonstrations which are the most difficult for the LLM to predict. In this paper, we implemented the algorithm based on the relevance generation approach. We define a score for each demonstration as the probability of the LLM generating the corresponding relevance label given a query and passage. The demonstrations with the lowest scores are applied.
$\bullet$ K-means: K-means is another static demonstration selection approach. This method clusters all the demonstrations in the pool into $k$ clusters and then selects $k$ demonstrations closest to each cluster center for ICL. We use the E5 (Wang et al., 2022) model to obtain the demonstration embeddings for clustering.
$\bullet$ BM25 (Robertson and Zaragoza, 2009): BM25 is a widely-used sparse retriever. We apply it to retrieve demonstrations that are most similar to the test query.
$\bullet$ SBERT (Reimers and Gurevych, 2019): We use Sentence-BERT as the off-the-shelf demonstration retriever following (Rubin et al., 2022) [^1] The checkpoint is from https://huggingface.co/sentence-transformers/paraphrase-mpnet-base-v2.. We use SBERT to encode all the demonstrations in the pool and retrieve the most similar ones.
$\bullet$ E5 (Wang et al., 2022): E5 is another off-the-shelf dense retriever. Following Wang et al. (2023), we use the same retrieval method as SBERT based on e5-base-v2 checkpoint [^2] https://huggingface.co/intfloat/e5-base-v2.
5.2. Main Results
We compare DemoRank with baselines in 3-shot ICL and the main results are shown in Table 1. From the results, we draw the following observations:
(1) Our framework DemoRank shows the best performance on all datasets and outperforms the second-best baselines with p-value $<$ 0.05 on most datasets, which indicates the significance of the improvements. For example, DemoRank outperforms the second-best model E5 on HotpotQA by about 4 points, and the second-best model BM25 on FEVER by about 8 points. The lack of significance on DL19 and DL20 is likely due to the small dataset size (only a few dozen queries), causing higher result variance.
(2) Random selection and rule-based methods (such as K-means and DBS) struggle to achieve performance improvements and may even be harmful (e.g., on the HotpotQA dataset). This is quite different from the phenomenon observed in the NLP field (Wang et al., 2023; Li et al., 2023), indicating that the demonstration selection for passage ranking is a very complex task.
(3) The similarity-based demonstration selection methods (e.g., BM25 and E5) perform well compared to other baselines. This suggests that the similarity between the demonstration and the input also helps select useful demonstrations. However, compared to DemoRank, they still fall significantly behind, proving that DemoRank can better model the relationships between the input and demonstrations, as well as the demonstration dependencies, thereby helping to select more effective few-shot demonstrations.
(4) The improvement from DemoRank compared to 0-shot varies greatly across datasets. For instance, DemoRank yields a 1-2 point gain on MSMARCO passage ranking datasets (MSMARCO, DL19, DL20) but nearly a 7-point boost on HotpotQA. We speculate that this is due to the varying difficulty levels of different datasets. HotpotQA is a multi-hop QA dataset, which requires more complex query understanding and reasoning than MSMARCO passage ranking datasets. Therefore, introducing demonstrations can better help the LLM understand the task, leading to greater improvements.
5.3. Analysis
In this section, we discuss different variants of DemoRank, compare DemoRank with different supervised passage rankers, and evaluate its performance using different demonstration numbers, unseen datasets, different LLM passage rankers and different demonstration retrieval numbers.
5.3.1. Different Variants of DemoRank
To understand the effectiveness of each component in DemoRank, we further evaluate different variants of DemoRank. We conduct the experiments on FEVER, NQ and DL19 with 3-shot ICL, shown in Table 2. First, we remove our demonstration reranker DReranker from DemoRank and only use demonstrations retrieved by our demonstration retriever DRetriever, denoted as “- DemoRank w/o DReranker”. We can see that removing DReranker causes about 1.5 points drop, which proves the effectiveness of our DReranker. Secondly, to further validate the effectiveness of our dependency-aware reranking approach, we introduce another reranker variant that ignores the demonstration dependency, denoted as “DemoRank w/o Dependency”. Specifically, given the demonstrations $Z_{r}$ retrieved by our DRetriever, this variant uses LLM to calculate a score of each candidate without considering other candidates, following the same scoring strategy as DRetriever. After that, the ranking list of scored demonstrations is used to fine-tune a demonstration reranker with the pairwise loss RankNet. Without considering the demonstration dependency, this variant lags behind DemoRank by about 2 points on FEVER and 1 point on average, which further proves that our dependency-aware reranking approach helps select more effective few-shot demonstrations.
We also investigate the effectiveness of ranking loss $L_{r}$ and the in-batch negatives on DRetriever training. Note that “DemoRank w/o DReranker” represents only applying DRetriever. Removing $L_{r}$ (denoted as DRetriever w/o $L_{r}$) causes about 0.6 points drop on average, which proving the effectiveness of ranking signals in scored demonstrations. Besides, incorporating the in-batch negatives into contrastive learning (denoted as DRetriever w/o IBN) brings no significant improvement and even causes a performance drop on DL19. This is because the utility of demonstrations in ranking tasks is not directly related to their similarity with the training input and the randomly sampled in-batch demonstrations may have a high score and act as a positive. Considering the lack of performance improvement and the additional computational overhead, we did not use in-batch negatives.
Lastly, we also validate the training effectiveness of our DReranker on different demonstration retrievers by replacing our trained DRetriever with E5 in our whole framework, denoted as $\text{DemoRank}_{\text{E5}}$. From the results, we can see that $\text{DemoRank}_{\text{E5}}$ significantly improves E5, which proves that our DReranker’s training is flexible and not restricted by specific demonstration retriever.
Table 2. Results (NDCG@10) of different variants.
Method FEVER NQ DL19 Average